Use asymptotes and a sign-analysis to sketch the function, ℎ(?).Use interval notation to denote theintervals of positive and negative derivatives, and positive and negative concavity.Expressall critical points and points of inflectionusing derivative notation.Label each axiswith the correct variables

h(x)= (2x^2) / (x^2-9)
Answer #1

I wen. Here f(x) = (x = q) 4 x 2 - 473_36x - 4x = (²9)2 - H=-3,3, - 362 J (229)2 (212294 but f(x)= 0 x=0 Here f(x) is not defезик , - -- - +2 > y = 2 X=-3 tox=3 Hasus The vertical Asymtotes of f(x) at x=-3 and x=3. Horizontal Alym totes of f(x) at y